OpenStudy (anonymous):

Determine whether the sequence converges or diverges. if it converges, find the limit. a sub n= (7+7n)/(9+3n)

OpenStudy (turingtest):

just take the limit as \(n\to\infty\)

OpenStudy (anonymous):

I really dont know how to approach these types of problems. I am trying to learn just from a book. And as you may know, calculus is hard to learn from a book

OpenStudy (turingtest):

we want to know if the sequence approaches a finite value as n goes to infinity, so just take the limit if it exists, it means that that is what the sequence converges to

OpenStudy (anonymous):

it converges.....and limit is 7/3

OpenStudy (sburchette):

L'Hopital's rule might help

OpenStudy (anonymous):

so you take the derivative so it would be 7/3 and thats how you get the answer

OpenStudy (sburchette):

If you use L'Hopital's rule, you will take the derivative of the top and bottom separately.

OpenStudy (turingtest):

l'Hospital is unnecessary; you can just divide top and bottom by n

OpenStudy (anonymous):

why do you divide by n?

OpenStudy (anonymous):

Or just observe. Neglect the terms which dont approach infinity. Take ratio of the coefficient of n. Although it's the same thing, really.

OpenStudy (turingtest):

\[\lim_{n\to\infty}{7+7n\over9+3n}\]if we try to take the limit now we get\[\frac\infty\infty\]but if we divide top and bottom by n first we get\[\lim_{n\to\infty}{\frac7n+7\over\frac9n+3}={\frac7\infty+7\over\frac9\infty+3}={0+7\over0+3}=\frac73\]

OpenStudy (anonymous):

do you always approach infinity?

OpenStudy (turingtest):

yes because that is what they are asking by asking if \(\{a_n\}\) converges they are asking of \(\{a_n\}\) approaches some finite value as \(n\to\infty\) that is, do the terms of the sequence ever "settle down" to any real number

OpenStudy (anonymous):

oh ok thank you!